OpenStudy (anonymous):

please i need your help Application of counting Two cards are drawn at random from an ordinary deck of 52 cards (no jokers). Find the probability that the 2-card hand contains the given card. b.No less than 1 diamond @tkhunny @satellite73

OpenStudy (anonymous):

no less than one diamond means one or two

OpenStudy (tkhunny):

Really big hint: p(no less than 1 diamond) = 1 - p(no diamond)

OpenStudy (anonymous):

compute the probability you get no diamonds, subtract it from one

OpenStudy (anonymous):

\(\frac{39}{52}\times \frac{38}{51}\) should be the probability you get no diamonds, i believe

OpenStudy (anonymous):

is it 39/52 im seeing

OpenStudy (anonymous):

what if im asked to find the probabitly of b. no deuces(2's) @satellite73